if a car is travelling at 60mph, how much meters will it travel in 2.8 seconds!?
Answer by srinivas.g(540)

speed =60 miles per hour
=60 * 1609.34 meter hour (1 mile =1609.34 meters)
= 96560.4 meters per hour
= 96560.4 /60 meters per minute ( 1 hr =60 min)
=1609.34 meters per min
= 1609.34 /60 meters per second ( 1 min =60 seconds)
=26.8223 meter per second
distance traveled in 2.8 seconds = 26.8223*2.8
= 75.1025 meters
Result :75.1025 meters